the sum of two numbers is 63. the first number is 2/3 of the second number. what is the number
Answer by algebrahouse.com(1659)

x = second number
(2/3)x = first number {first number is 2/3 of the second}

x + (2/3)x = 63 {sum of the two numbers is 63}
3x + 2x = 189 {multiplied entire equation by 3 to eliminate fraction}
5x = 189 {combined like terms}
x = 189/5 {divided each side by 5}
(2/3)x = 126/5 {substituted 189/5, in for x, into (2/3)x

first number is 126/5 or 25 1/5
second number is 189/5 or 37 4/5

For more help from me, visit: www.algebrahouse.com